OpenStudy (babyslapmafro):

Find the volume of the solid that results when the region above the x-axis and below the ellipse is revolved about the x-axis (click to see equation).

OpenStudy (babyslapmafro):

\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (anonymous):

try solving for y, then find the intercepts, then integrate y using those intercepts as limits of integration